### v1

# The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profiles

Abstract
We examine the splashback structure of galaxy clusters using hydrodynamical simulations from the GIZMO run of The Three Hundred Project, focusing on the relationship between the stellar and dark matter components. We dynamically decompose clusters into orbiting and infalling material and fit their density profiles. We find that the truncation radius $r_{\mathrm{t}}$, associated with the splashback feature, coincides for stars and dark matter, but the stellar profile exhibits a systematically steeper decline. Both components follow a consistent $r_{\mathrm{t}}{-}\Gamma$ relation, where $\Gamma$ is the mass accretion rate, which suggests that stellar profiles can be used to infer recent cluster mass growth. We also find that the normalisation of the density profile of infalling material correlates with $\Gamma$, and that stellar and dark matter scale radii coincide when measured non-parametrically. By fitting stellar profiles in projection, we show that $r_{\mathrm{t}}$ can, in principle, be recovered observationally, with a typical scatter of $\sim 0.3\,R_{200\mathrm{m}}$. Our results demonstrate that the splashback feature in the stellar component provides a viable proxy for the cluster’s physical boundary and recent growth by mass accretion, offering a complementary observable tracer to satellite galaxies and weak lensing.
keywords:
galaxies: clusters: general – galaxies: halos – cosmology: theory – large-scale structure of the universe - methods: numerical
pubyear: 2025pagerange: The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profilesReferences
1 Introduction
The precise definition of a galaxy cluster’s outer boundary is fundamental to studies of cluster formation, evolution, and cosmology (e.g. Kravtsov & Borgani, 2012; Morandi & Sun, 2016; Haggar et al., 2024). Traditional definitions — deduced from friends-of-friends (FoF) linking lengths or spherical overdensity radii —play a key role in determining cluster masses and membership (e.g. Eke et al., 2004; Lukić et al., 2009), yet these are choices driven primarily by what is practicable with data rather than what is most physically meaningful (e.g. White, 2001; Old et al., 2014, 2015). Establishing a boundary that is physically meaningful and practicably measurable via observations remains a major challenge.
A promising physically motivated boundary is provided by the splashback radius, which marks the apocentre of recently accreted material orbiting within the cluster potential (e.g. Diemer & Kravtsov 2014; Adhikari et al. 2014; More et al. 2015; Diemer 2022). This radius delineates the transition between orbiting and infalling matter, and manifests as a steepening in the halo density profile (which, we note, has historically—and erroneously—been conflated with the splashback radius itself). The radius of this steepest slope is of particular interest because there is now considerable evidence that this steepening has been measured in observations of real galaxy clusters (e.g. More et al. 2016; Baxter et al. 2017; Nishizawa et al. 2018; Chang et al. 2018a; Contigiani et al. 2019; Xu et al. 2024). This is exciting because there is good reason to expect that splashback radius offers a direct probe of halo assembly history and mass accretion rates, thereby connecting cluster outskirts to cosmological growth (More et al., 2015; Deason et al., 2021; Dacunha et al., 2025).
Observational detections of splashback features have been reported using satellite galaxy counts (e.g. More et al. 2016; Baxter et al. 2017; Nishizawa et al. 2018), weak gravitational lensing (e.g. Chang et al. 2018a; Contigiani et al. 2019; Xu et al. 2024), and the Sunyaev-Zel’dovich (SZ) effect (e.g. Shin et al., 2019; Zürcher & More, 2019; Adhikari et al., 2021) However, these come with their own unique limitations. For example, in the case of satellite galaxy counts, accurate measurement of splashback features depends on the properties of satellite tracers in a given cluster (e.g. O’Neil et al., 2022), and accounting for the complexity of their kinematics (e.g. Aung et al., 2021) and structure of the local cosmic web (e.g. Lebeau et al., 2024; Sun et al., 2025). Satellite galaxies are also biased tracers of the cluster potential, with the most massive tending to accumulate more centrally due to dynamical friction (Ludlow et al., 2009; Adhikari et al., 2016). In the case of weak lensing, individual systems may have only a small number of lensed objects.
A promising alternative probe is the diffuse intra-cluster light (ICL) that is formed from the remnants of disrupted galaxies. This material is a potentially powerful tracer of cluster outskirts (Montes & Trujillo, 2019; Contini, 2021), and work by Deason et al. (2021) has shown a direct correspondence between the caustics in the total stellar and dark matter density, indicating a dynamical link between these components. However, there remain important questions about the detailed nature of the stellar splashback structure, its relation to the underlying dark matter, and its observational accessibility, which we seek to answer in this paper.
A deeper understanding of the stellar splashback feature is crucial for a number reasons. The distribution of stars in a cluster reflect physical processes that affect satellite galaxies and their dark matter halos (such as tidal stripping and dynamical friction) and so their distribution encodes information about the interplay between baryons and dark matter in the cluster outskirts (e.g. Brown et al., 2024), which, when combined with splashback and shock structures in the stars and gas (e.g. Deason et al., 2021; Towler et al., 2024; Zhang et al., 2025) can be used to infer dynamical history. Similarly, projected stellar density profiles provide a complementary means of constraining cluster growth histories (e.g. Caminha et al., 2017; Montes & Trujillo, 2018; Kluge et al., 2025; Joshi et al., 2025). This is especially valuable when supplemented by weak lensing and SZ effect measurements; combining different datasets, affected by projection in different way, allow for a more robust inference of splashback features (e.g. Adhikari et al., 2021).
In this paper, we draw upon the statistical sample of massive galaxy clusters from The Three Hundred collaboration’s hydrodynamical galaxy formation simulation suite (cf. Cui et al., 2018, 2022) to study the relationship between splashback features in the dark matter and stellar components. This is a mass complete sample of clusters drawn from a 1 $h^{-1}$ Gpc box, which have a diversity of assembly histories and larger-scale environments. These simulations self-consistently model the physics of galaxy formation, including star formation, black hole growth, supernovae, and AGN feedback. By investigating the stellar splashback structure and its relation to dark matter in hydrodynamical simulations, our aim is to establish a physically grounded framework for interpreting stellar density profiles as tracers of cluster assembly. This approach promises to enhance the utility of stellar observables—such as the ICL—in constraining cluster boundaries and mass accretion histories, thereby broadening the toolkit available for cluster cosmology.
The remainder of this paper is organised as follows. §2 summarises the simulation data used in our analysis, and §3 describes our approach to decomposing and characterising the galaxy cluster structure into orbiting and infalling components. In §4 we present our key results, and we discuss these results in the context of previous studies in §5. Finally, we summarise our key results and identify future directions in §6
2 The data
For this work we use data from the Three Hundred Project, a suite of hydrodynamical re-simulations of the most massive 324 galaxy cluster halos in the dark matter-only MultiDark-Planck 2 (MDPL2) simulation (cf. Klypin et al., 2016), a $1h^{-1}$ Gpc box on a side. These clusters have virial masses in the range $6.4\times 10^{14}h^{-1}M_{\odot}\lesssim M_{200}\lesssim 2.6\times 10^{15}h^{-1}M_{\odot}$, where $M_{200}$ is the mass corresponding to an overdensity criterion of 200 times the critical density at that epoch. Each resimulated region extends 15 $h^{-1}$Mpc from the cluster centre at zz=0, corresponding to several virial radii.
In this work we use hydrodynamical simulations run with the galaxy formation model simba (cf. Davé et al., 2019), implemented in the gizmo code (cf. Hopkins, 2015). The gravity solver used in the model derives from gadget (cf. Springel, 2005). gizmo uses a moving mesh scheme to solve hydrodynamics. This overcomes the limitations of classic SPH and captures fluid mixing and instabilities. Galaxy formation processes (e.g. radiative cooling, star formation and feedback, black hole formation and growth, multiple modes of black hole feedback) are modeled using a variant of simba (cf. Davé et al., 2019) that has been calibrated for cluster scales (cf. Cui et al., 2022). gizmo was run on regions from both the $3840^{3}$- and $7680^{3}$-particle MDPL2 boxes, which we refer to as GIZMO_3k and GIZMO_7k, respectively. In this work we mainly present results from the high resolution run, but we verify that the lower resolution run gives identical results.
[…]
Data are saved in 129 snapshots from $z=16.98$ to $z=0$. For each cluster, we use group catalogues constructed with the AHF halo finder (cf. Knollmann & Knebe, 2009), which includes information about the stellar and gas content of the main halo and its substructures.
3 Methods
In this section we outline the algorithm used to decompose the clusters into their orbiting and infalling components. We then go over the calculation and fitting of the density profiles.
3.1 Orbiting-infalling decomposition
To identify the splashback of stars and dark matter in the simulated clusters, we classify particles as “orbiting” or “infalling” depending on if the particle has passed through at least one pericenter about the center of mass of the cluster. This is done by recording a change in the sign of a particle’s physical radial velocity (peculiar plus Hubble flow) from negative to positive. To filter out radial velocity sign flips due to orbits within subhalos, we also require that particles have traversed an angle of at least $\pi/2$ around the halo center since entering the cluster region. Any particle that has not had such a sign change is regarded as “infalling”. Though this classification technically considers positive radial velocity particles beyond the turn-around radius as “infalling”, we limit our analysis of the density profiles to radii safely within this boundary.
The result of the orbiting-infalling decomposition for the most massive cluster in the GIZMO_3k run is shown in Figure 1. The top row shows the $x$-$y$ projection of the dark matter (left) and stars (right), while the bottom row is their radial phase space (with Hubble flow added). The phase space shows a clear separation between the infalling stream and the orbiting material, with the infalling material abruptly transitioning into orbiting material at the point of pericenter, $v_{\rm r}=0$. The orbiting particles at the largest radii still have some positive radial velocity, meaning they are still moving outwards and have not quite “splashed back”. Visually, the distribution of the dark matter is well traced by the stellar distribution, and both the orbiting dark matter and orbiting stars extend out slightly beyond the virial radius $R_{\rm vir}$ (Bryan & Norman 1998 definition). See e.g. Diemer (2017); Bakels et al. (2021) for discussion on the extent of the orbits of halo material.
Figure 1: Cluster 001 from the GIZMO_3k run, decomposed into orbiting (coloured) and infalling (white) particles. The top row shows the $x$-$y$ projection of the dark matter (left) and stars (right), while the bottom row shows their radial phase space. The virial radius (see Bryan & Norman 1998 for definition) is marked.
Figure 2: The density profiles of the orbiting and infalling components, fit using models 1 - 3. The BCG component is confined to the inner regions, while the stellar “halo” takes on a similar shape to the orbiting profile of the dark matter. The logarithmic slope profiles are shown in the bottom panels.
3.2 Cluster density profiles
The density profiles of the clusters are calculated at $z=0$ in logarithmically-spaced bins between $0.01\,{\rm Mpc}$ and $15\,{\rm Mpc}$ from the position of peak density, as determined by AHF. We further bin the radial shells into segments of equal solid angle and compute the median, which suppresses local fluctuations in density due to anisotropic substructure. For the lower-resolution GIZMO_3k run, we choose 128 radial bins and 48 angular bins for the dark matter (average of $\sim$4,000 particles per radial bin), and 64 radial bins and 12 angular bins for the stars (average of $\sim$600 particles per radial bin). For the high-resolution GIZMO_7k run, we choose 128 radial bins and 48 angular bins for both the dark matter and the stars (average of $\sim$30,000 and $\sim$3,000 particles per radial bin, respectively). As well as the total density profile, we also calculate the density profiles of the orbiting and infalling material using the same binning. To calculate the logarithmic density slope profiles, we take the derivative using a $4^{\rm th}$-order Savitzsky-Golay filter over the 20 nearest radial bins for the 128 bin profiles, and the nearest 10 radial bins for the 64 bin profiles. The profiles are calculated at $z=0$ using all particles of the relevant type, whether gravitationally bound or not.
The dark matter and stellar density profile of the most massive cluster in the GIZMO_3k run (cluster 001) is shown in Figure 2. The profile of the orbiting material is shown in red, while that of the infalling material is in grey. The bottom panels show the corresponding logarithmic slope profiles. The radius is normalized by $R_{200{\rm m}}$, the radius enclosing an average density equal to $200$ times the mean matter density of the universe, $\rho_{\rm m}$. All overdensity radii and masses we use in this work are calculated using the total density (stars, gas, and dark matter). The outer profiles of the dark matter and stars ($\gtrsim 0.1\times R_{200{\rm m}}$) are—at least qualitatively—very similar in shape (though very different in magnitude): both feature a rapidly steepening orbiting profile and an infalling profile that decreases much more gradually with radius, eventually overtaking the orbiting density at $\sim R_{200{\rm m}}$ before plateauing to the background density. In the inner regions the profiles are very different, with the stellar profiles having an enhancement in the density due to the cluster’s central galaxy (the “brightest cluster galaxy”, or BCG). The logarithmic slope of the total profiles exhibit the well-known minimum near $R_{200{\rm m}}$, which occur at the same radius for the dark matter and stars and coincide with the truncation in the slope of the orbiting profile.
3.3 Profile fitting functions
To understand how the stellar orbiting and infalling profiles quantitatively relate to those of the dark matter we use the fitting functions introduced by Diemer (2023), which have been shown to recover the orbiting and infalling profiles of the dark matter from the total profile reasonably well (Diemer, 2025). The fitting function for the orbiting profile is an Einasto profile that is exponentially suppressed by a truncation term:
$\rho_{\rm orb,DM}(r)=\rho_{\rm s}e^{S(r)},$
where
$S(r)=-\frac{2}{\alpha}\left[\left(\frac{r}{r_{\rm s}}\right)^{\alpha}-1\right]-\frac{1}{\beta}\left[\left(\frac{r}{r_{\rm t}}\right)^{\beta}-\left(\frac{r_{\rm s}}{r_{\rm t}}\right)^{\beta}\right].$ (1)
Here $r_{\rm s}$ and $\rho_{\rm s}=\rho(r_{\rm s})$ are the scale radius and scale density, $\alpha$ is the exponent of the power-law inner slope, and $r_{\rm t}$ and $\beta$ are the radius and steepness of the orbiting profile truncation.
Though this function describes the orbiting dark matter well, it is not well-suited to fitting the inner part of the stellar profile where the BCG becomes important. We therefore add a Sérsic “BCG” term to Equation 1. We refer to the part described by Equation 1 as the stellar “halo”, so that the orbiting fitting function for the stars is given by
$\rho_{\rm orb,\ast}(r)$ $=\rho_{\rm halo}(r)+\rho_{\rm BCG}(r)$
$=\rho_{\rm s}e^{S(r)}+\rho_{0}\exp\left\{-c\left[\left(\frac{r}{r_{0}}\right)^{k}-1\right]\right\}$ (2)
where $c$ and $k$ control the shape of the BCG profile and we refer to $r_{0}$ and $\rho_{0}=\rho(r_{0})$ as the BCG radius and characteristic density, respectively. We do not assume any relationship between $c$ and $k$ and instead allow them to vary freely when fitting. We also emphasise that although $\rho_{\rm orb,DM}$ and $\rho_{\rm halo}$ have the same functional form, we refer to them with different names because they represent different quantities—namely the dark matter orbiting profile and the “halo” part of the stellar orbiting profile.
Figure 3: Comparison between $r_{\rm t}$ and $\beta$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The cyan squares are from fitting the $\Gamma$-binned median profiles.
The infalling model is given by the following function that flattens out to fixed densities at small and large radii:
$\rho_{\rm inf}(r)=\rho_{\infty}\left(\frac{\delta_{1}}{\sqrt{(\delta_{1}/\delta_{\rm max})^{2}+(r/R_{200{\rm m}})^{2s}}}+1\right),$ (3)
where $\delta_{1}$ is the normalization, $s$ controls the slope, and $\rho_{\infty}$ is the density at large radii. Although the statistical contribution from large-scale structure begins to dominate the density profile at large radii, we only fit out to a few times $R_{200{\rm m}}$ and so we do not encounter this transition. Since the shape of the stellar infalling profile is similar to that of the dark matter, we expect it to be well-fit by Equation 3.
We fit the profiles using least-squares minimization in the logarithm of the density. The radial bins are weighted equally and the fitting is done between $4\,R_{200{\rm m}}$ and the convergence radius predicted by Power et al. (2003) and Ludlow et al. (2019) for the dark matter-only case. The results of fitting the models 1 - 3 to the orbiting and infalling profiles of an individual cluster (cluster 001) is shown in Figure 2.
As well as fitting the orbiting and infalling profiles of the individual clusters, we also fit profiles averaged in bins of accretion rate to investigate any possible systematic differences. We define the accretion rate at $z=0$ as
$\Gamma\equiv\Gamma_{200{\rm m}}=\frac{\Delta\log\,M_{200{\rm m}}}{\Delta\log\,a},$ (4)
where $a$ is the scale factor and $\Delta$ represents a change over one dynamical time, defined as the crossing time $2\,R_{200{\rm m}}/V_{200{\rm m}}$. At $z=0$, one dynamical time in the past corresponds to $a\approx 0.769$. Here, as with everywhere else in this work, $M_{200{\rm m}}$ is calculated using the total density (stars, gas, and dark matter). Progenitor halos are identified as those that contain the largest fraction of the most-bound particles of the $z=0$ halos. We choose six equal-width bins between $\Gamma=0$ and $\Gamma=6$ and calculate the median logarithmic density profile in each bin, in units of $r/R_{200{\rm m}}$. The first four bins contain between 51 and 102 clusters, while the remaining two contain 17 and 10.
4 Results
In this section, we compare the fit parameters from the stellar and dark matter profiles, focusing specifically on the parameters $r_{\rm t}$ and $\beta$ that describe the splashback region. The first subsection shows results from fitting the 3D orbiting and infalling profiles, and so represent the “true” parameters. In the subsequent subsection, we present results from fitting the projected total profiles to determine how well the truncation parameters can be recovered from observed profiles. We leave the comparison of the infalling component to Section 4.3 and discussion of the inner profile to Section 4.4.
4.1 Fit parameter relations
Figure 3 shows how $r_{\rm t}$ (left) and $\beta$ (right) compare between dark matter and stars. $r_{\rm t}$ follows a one-to-one relation with relatively small scatter—in other words, the splashback radius of the stars and dark matter in clusters is the same on average. This confirms what was strongly hinted at by the Deason et al. (2021) results on the radius of steepest slope. On the other hand, $\beta$ does not follow a one-to-one relation, with the truncation in the stars generally being steeper than that of the dark matter. This is also consistent with the Deason et al. (2021) finding that the steepest slope is steeper for the stars than the dark matter. The points are colored by accretion rate (Equation 4), which reveals a clear gradient in $r_{\rm t}$ and a less clear but still noticeable gradient in $\beta$, with each increasing with decreasing $\Gamma$. That is, the splashback boundary is sharper and is found at a larger fraction of $R_{200{\rm m}}$ for slowly accreting clusters. The cyan squares show the $r_{\rm t}$ and $\beta$ from fitting to the median of the $\Gamma$-binned profiles, and they are consistent with the individual measurements.
Figure 3 shows that the outer orbiting profile—i.e. the splashback region—is very similar between dark matter and stars, albeit with the stellar density dropping off more steeply on average. We should therefore expect the splashback region of the stars in clusters to be an accurate tracer for the splashback region of the dark matter.
Figure 4: Comparison between the radius of steepest slope, $r_{\rm steep}$ and $r_{\rm t}$ for the dark matter and stars. The points are colored by accretion rate (Equation 4). The cyan squares are from $\Gamma$-binned median profiles. The black dashed lines are linear fits given by $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.41$ for the dark matter and $r_{\rm steep}/R_{200{\rm m}}=0.83(r_{\rm t}/R_{200{\rm m}})+0.37$ for the stars.
Figure 5: $r_{\rm steep}$ and $r_{\rm t}$ as a function of accretion rate (Equation 4). The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), while in the bottom row it is the same relation rescaled using the best-fit $r_{\rm t}$-$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025).
Figure 6: The decomposed stellar density profile of the same cluster from the GIZMO_7k run (cluster 005) in 3D (left) and in projection (right), along with the best-fit curves. The projected density has been normalized by the mean matter density multiplied by 30 Mpc, the diameter of the cluster region.
The truncation radius is a more direct measure of the size of the splashback boundary (technically the orbiting boundary, since the outermost particles may not have reached apocenter yet) than the radius of steepest slope, which is determined in part by the infalling material. In perfectly spherical collapse, the truncation of the orbiting profile is infinitely sharp and the two radii coincide. However, in realistic simulations, mass is accreted anisotropically and particle orbits can take on a variety of shapes, which has the effect of smoothing out the edge of the orbiting profile. In this case, the radius of steepest slope is not simply related to the edge of the orbiting material, but is rather the point at which the infalling profile becomes dominant enough to strongly influence the total slope. In Figure 4 we plot $r_{\rm steep}$ measured from the total profiles against $r_{\rm t}$. Again, the points are coloured by $\Gamma$ and show the gradient expected from the now well-known $r_{\rm steep}$-$\Gamma$ relation. The cyan points are measurements of $r_{\rm steep}$ from the $\Gamma$-binned median total profile. The $r_{\rm t}$-$r_{\rm steep}$ relation seems to be described well by a line with slope of $\sim 0.83$ and $r_{\rm steep}(r_{\rm t}=0)/R_{200{\rm m}}\approx 0.41$ for the dark matter and $0.37$ for the stars (black dashed lines). On average, the radius of steepest slope is $\sim 0.25\,R_{200{\rm m}}$ larger than the truncation radius.
The gradient with accretion rate seen in the radii in Figures 3 and 4 reflect the now well-established relationship between the splashback radius and accretion rate. This relationship was first reported by Diemer & Kravtsov (2014) for the radius of steepest slope in dark matter-only cosmological simulations, while its theoretical basis was established by Adhikari et al. (2014) using a model based on spherical collapse. The relation was shown by More et al. (2015) to be well-described by an exponential of the form $r_{\rm steep}/R_{200{\rm m}}=a+be^{-\Gamma/c}$, and this same function was shown by Deason et al. (2021) to hold for the stellar $r_{\rm steep}$ of 34 C-EAGLE clusters. Recently, Diemer (2025) showed that the relationship between $r_{\rm t}$ and $\Gamma$ for stacked halos is also well-described by an exponential that has remarkably little dependence on halo mass or cosmology.
Figure 7: Comparison between 3D and projected $r_{\rm t}$ and $\beta$. The points are colored by accretion rate (Equation 4). The cyan squares are from the $\Gamma$-binned median profiles. The black dashed line is a linear fit.
In Figure 5 we plot the accretion rate dependence of the stellar $r_{\rm t}$ and $r_{\rm steep}$. For comparison, we also show the corresponding plots for the dark matter. The light blue points are individual clusters, while the dark blue points show the median in the same accretion rate bins as we use to calculate median profiles. The measurements from the median profiles are shown in red squares. The increased transparency on the right-most points indicates that they correspond to bins with <20<20 clusters. The fitting function for $\Gamma_{200{\rm m}}-r_{\rm steep}$ from More et al. (2015) is plotted on the top row and agrees well with our data. The fitting function for $\Gamma_{200{\rm m}}-r_{\rm t}$ from Diemer (2025) is plotted on the bottom row and agrees at large $\Gamma_{200{\rm m}}$, but rises more rapidly than our data towards the small end. For comparison, we plot the More et al. (2015) relation rescaled according to the $r_{\rm t}$-$r_{\rm steep}$ relation we find from Figure 4, and it seems to be a better fit.
Comparing the dark matter and the stars, we see that they follow the same relationship with $\Gamma_{200{\rm m}}$ and that the scatter is very similar. The accretion rate can therefore be estimated from measurements of the stellar $r_{\rm t}$—assuming one has access to the 3D orbiting profile of the stars. However, observations only have access to the total stellar density, and in projection at that. We therefore now turn to the projected total profiles to assess how these results can be applied to observations.
4.2 Recovering $r_{\rm t}$ from observable profiles
So far, we have shown results from fitting the three-dimensional orbiting and infalling profiles. However, real clusters are observed in projection and information about the orbiting and infalling material is not directly accessible. In this section, we show results from fitting the total, projected density profiles of the stars, as well as those of the dark matter for comparison. This means using no information about the orbiting and infalling identification and instead fitting the density profile with the total model, which in projection we write as $\Sigma(r)=\Sigma_{\rm orb}(r)+\Sigma_{\rm inf}(r)$.
We calculate the projected density profiles using the same radial binning as we used for the 3D profiles. The angular binning is also done with the same number of bins, but this time in equal angle rather than equal solid angle. We project the entire cluster region along the $x$, $y$ and $z$ axes, and take the mean of the resulting profiles. While choosing different windows along the line of sight will produce distinct profiles, we mainly concern ourselves with the recoverability of the orbiting parameters rather than their actual values in projection. The calculation of the density slope profiles is the same as for the 3D case. For clarity, we refer to the projected density with the symbol $\Sigma$ and other quantities with upper case, e.g. $R_{\rm t}$ and $B$ for the truncation radius and steepness.
In projection, the shape of the density profile is modified. For the stars, the splashback region and BCG-stellar halo transition are smoothed out. This can be seen in Figure 6, which compares the 3D and projected stellar profile of one of the clusters in the GIZMO_7k run. Despite this, the profiles are still well-fit by the functions 1 - 3, generalizing earlier work on the similarity between 3D and projected total profiles (e.g. Merritt et al. 2005). It is reasonable to expect that the parameters of the 3D and projected profiles should be related. Figure 7 compares the truncation radius and truncation steepness from fitting the 3D and projected orbiting profiles (not the total profiles), for the dark matter (left) and stars (right). The projected $R_{\rm t}$ is slightly smaller than the 3D $r_{\rm t}$ by a factor of $\sim 0.92$ for the dark matter and $0.94$ for the stars (black dashed lines). This is similar to the factor of $0.9$ found by Deason et al. (2021) when comparing the 3D and projected steepest slope radii. The projected $B$ of the dark matter is also smaller than its 3D counterpart by a factor of $\sim 0.87$. This may also be the case for the stars, but the scatter is significantly larger.
We have chosen to show the results from fitting the orbiting profiles in Figure 7 to see the “true” effect of projection on the outer profile. Fitting the total profile, as is required when information about the orbiting and infalling material is inaccessible, inevitably sacrifices some accuracy as the model attempts to recover the “best-fitting” orbiting and infalling profiles. As shown by Diemer (2025), the model is able to do this relatively successfully—at least for binned profiles. In Figure 8, we compare the $R_{\rm t}$ obtained from fitting the orbiting profile (which we refer to as “component”) to that obtained from fitting the total profile, where both are in projection. We find that fitting the total profiles introduces a scatter of $\sim 0.3$ around the “true” $R_{\rm t}$, for both the dark matter and the stars. The median total $R_{\rm t}$, calculated in five bins in the component $R_{\rm t}$ and denoted by the red circles, shows that there is no significant $R_{\rm t}$-dependent systematic error introduced by fitting the total profiles of the dark matter, while there is possibly a slight systematic error for the stars.
Figure 8: Comparison between $r_{\rm t}$ obtained by fitting the orbiting profile (component) and the total profile. The points are colored by accretion rate (Equation 4). The cyan squares are from the $\Gamma$-binned median profiles.
Figure 9: $R_{\rm steep}$ and $R_{\rm t}$ as a function of accretion rate (Equation 4). The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), rescaled using the factor of $r_{\rm steep}/R_{\rm steep}=0.9$ found by Deason et al. (2021). The black dashed in the bottom row is the same curve further rescaled using the best-fit $r_{\rm t}$-$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025), rescaled by the slope of the best-fit line shown in Figure 7.
Figure 10: Best-fit parameters for the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The green and red squares are the parameters of the mean and median (respectively) $\Gamma$-binned profiles.
Figure 9 is the equivalent of Figure 5 for the $R_{\rm t}$ and $R_{\rm steep}$ measured from the projected total profiles. This time, the More et al. (2015) relation for $r_{\rm steep}$ has been multiplied by the $r_{\rm steep}/R_{\rm steep}$ factor found by Deason et al. (2021), while the $r_{\rm t}$ relations have been multiplied by the $r_{\rm t}/R_{\rm t}$ factor we found from Figure 7.
The scatter in both $R_{\rm steep}$ and $R_{\rm t}$ is larger compared to the 3D and 3D component measurements, however this increase is more noticeable for $R_{\rm t}$ due to the error introduced when fitting to the total profiles (see Figure 8). Despite this, the accretion rate dependence is still discernable.
The $R_{\rm steep}$-$\Gamma$ plots feature a small group of points that lie below the relation at low $\Gamma$. These are cases where the $2^{\rm nd}$ caustic—an additional, usually shallower caustic at smaller radius—actually becomes steeper than the $1^{\rm st}$ caustic when viewed in projection. The measurement of $R_{\rm t}$ seems to be unaffected by this, suggesting the fitting method may have an advantage in these cases.
4.3 The infalling profile
So far we have focused on the outer orbiting profile of the stars and dark matter. In this section we turn to the parameters of the infalling profile—the profile of material that has yet to make its first pericentric passage—to see how they compare between the dark matter and stars.
Figure 10 shows the distribution of the parameters from fitting the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The green and red squares are the parameters of the mean and median (respectively) $\Gamma$-binned profiles.
Unsurprisingly, the asymptotic density $\rho_{\infty}/\rho_{\rm m}$ is much smaller for the stars, as seen in the top-left plot. The $\Gamma$-binned points all lie on top of one another, suggesting that $\rho_{\infty}/\rho_{\rm m}$ does not evolve with $\Gamma_{200{\rm m}}$. This makes sense since material far outside $R_{200{\rm m}}$ has not yet been accreted. By taking the mean of the median binned points, we find that the asymptotic dark matter density is $\rho_{\infty}^{\rm DM}/\rho_{\infty}^{\ast}\sim 182$ times that of the stars.
Figure 11: Best-fit infalling density at $r=0$ for dark matter (blue) and stars (red) in 3D and projection. The larger, dark points are from the median $\Gamma$-binned profiles. The bottom panels show the ratio between the blue and red points. The black dashed lines are lines of best fit to the large points.
The normalization of the stellar infalling profile, $\log_{10}\delta_{1}$, is smaller than its dark matter counterpart and increases with it, though not as a one-to-one relationship. The case is similar for $\log_{10}\delta_{\rm max}$, though the slope is closer to 1.
The infalling slope, $s$, is generally larger for the stars but has significantly larger scatter than the dark matter.
Based on the $\Gamma$-binned points, $\delta_{1}$, $\delta_{\rm max}$, and $s$ all appear to evolve with accretion rate. However, plotting these against accretion rate shows that, for $\delta_{1}$ and $s$, this is mostly scatter. $\delta_{\rm max}$, on the other hand, evolves unambiguously with accretion rate. The scatter in this relation is reduced when we instead consider the predicted infalling density at the center of the halo: $\rho_{\rm inf}(0)=\rho_{\infty}(\delta_{\rm max}+1)$. This is shown in Figure 11. The dependence on $\Gamma_{200{\rm m}}$ for both the stars (red) and dark matter (blue) is linear with minimal scatter. The slope of the stellar relation is slightly larger than that of the dark matter, both in 3D and projection. The ratio of the dark matter and stellar density is shown in green in the sub-plots.
4.4 The inner profile
In this section we turn to the inner density profile of the stars and discuss its relation to the dark matter.
In our model, the orbiting density profile of the stars is composed of a Sérsic BCG and a stellar “halo” fit with the truncated Einasto profile of Diemer (2025) (see Section 3.3). As with the unmodified Einasto profile, the shape of the inner profile is controlled by the inner slope exponent $\alpha$ and the scale radius $r_{\rm s}$—the radius at which the logarithmic slope equals 2-2. Figure 12 compares $r_{\rm s}$ and $\alpha$ between the stars and dark matter. Unlike for the outer profile parameters, the stellar inner profile parameters do not track the dark matter well. Though there is a population with $r_{\rm s}$ along the one-to-one line, the stellar $r_{\rm s}$ tends to underestimate that of the dark matter, while $\alpha$ has significant scatter around the one-to-one line.
Figure 12: Comparison between $r_{\rm s}$ and $\alpha$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The cyan squares show the results from fitting the $\Gamma$-binned median profiles.
Figure 13: Comparison between stellar and dark matter $\rho r^{2}$ profiles of the nine most massive clusters in the GIZMO_7k run. The stellar profiles have been rescaled by a fixed number $k_{\rm SHM}=138$. The sub-panels show the derivative of the profiles. The vertical gray lines mark the location of the peak of the dark matter profile.
The reason for this disagreement is the freedom in the inner profile allowed by the BCG term. The BCG is typically and order of magnitude more dense than the stellar halo (see Figures 2 and 6), meaning that, in log space, even a small change to the BCG fit can lead to a large change in the shape of the inner stellar halo profile.
To see if the disagreement between the stellar and dark matter $r_{\rm s}$ suggested by Figure 12 is real or merely an artifact of the fitting, we measure $r_{\rm s}$ non-parametrically by plotting the $\rho\,r^{2}$ profile, which peaks when the logarithmic slope of $\rho$ is 2-2. Figure 13 shows the $\rho\,r^{2}$ profiles of nine of the most massive clusters in the GIZMO_7k run, along with their logarithmic derivatives. The stellar profiles have been rescaled by a fixed factor of $k_{\rm SHM}=138$. The peaks in the stellar profile are more difficult to discern than those of the dark matter due to their proximity to the central density enhancement—nevertheless, in most cases they are very close to those of the dark matter, suggesting that the $r_{\rm s}$ of the stars should indeed coincide with that of the dark matter.
5 Discussion
5.1 Comparison to Previous Simulations
The agreement between the radius of steepest slope of the dark matter and stellar density in simulated galaxy clusters shown by Deason et al. (2021) suggests a close relationship between their splashback structure. We decomposed the stars and dark matter into their orbiting and infalling components and parameterized the edge of the orbiting profile by the truncation radius, $r_{\rm t}$, and slope, $\beta$ from Diemer (2025). We showed that the $r_{\rm t}$ of the stars and dark matter coincide, and that $\beta$ is generally larger for the stars. In other words, the splashback region is sharper for stars, but is at the same location as that of the dark matter.
This result is consistent with previous simulations. For example, Diemer (2017) demonstrated that the splashback radius is a robust halo boundary across a wide range of masses, accretion rates, and cosmologies, and provided fitting functions that match our results well. O’Neil et al. (2021) showed that baryonic physics has minimal impact on the location of the splashback radius, supporting our use of stellar profiles as reliable tracers. However, Lebeau et al. (2024) highlighted that the local cosmic web and associated accretion of galaxies from filaments can distort splashback measurements, especially in dynamically disturbed clusters like Virgo. These findings underscore the importance of accounting for environmental and dynamical diversity when interpreting splashback features.
The coincidence between the $r_{\rm t}$ means that the stellar $r_{\rm t}-\Gamma$ relation has the same form as that of the dark matter. This makes it a complimentary estimator of the accretion rate in clusters—assuming that $r_{\rm t}$ can be measured accurately. However, in observations, information of the orbiting and infalling components is inaccessible, and $r_{\rm t}$ must instead be estimated by fitting the total profile in projection. Given the freedom allowed by the full nine-parameter model of the total profile, the orbiting and infalling components may not always be accurately recovered.
When fitting the total profile, we find that the projected truncation radius, $R_{\rm t}$, has a standard deviation from the “true” $R_{\rm t}$ (measured from the orbiting profile) of $\sim 0.3\,R_{200{\rm m}}$ for both the dark matter and stars. There may also be a slight $R_{\rm t}$-dependent systematic difference in the case of the stars, which would produce a subtle change in the $R_{\rm t}$-$\Gamma$ relation, but this is unlikely to be discernable. While the exponential relationship can still be clearly seen, the scatter is larger than that of the $R_{\rm steep}$-$\Gamma$ relationship and so it is likely to be a poorer estimator of the accretion rate in observations.
5.2 Comparison to Observations
Gravitational lensing, Sunyaev–Zel’dovich (SZ) observations, and measurements of the distribution of cluster galaxies in projection provide complementary approaches to identifying the splashback boundary in galaxy clusters, each offering independent insight into the halo’s outer structure. Weak lensing directly probes the total matter distribution and has revealed a steepening in the density profile consistent with theoretical predictions of the splashback feature (e.g. Chang et al., 2018b; Contigiani et al., 2019). However, lensing measurements are inherently noisy at large radii due to shape noise and projection effects, and often require stacking to obtain robust profiles.
Chang et al. (2018b) conducted one of the earliest joint analyses of galaxy density and weak lensing profiles using DES Year 1 data, detecting a splashback-like steepening in both tracers. They measured a splashback radius of $r_{\mathrm{sp}}=1.34\pm 0.21\,h^{-1}\,\mathrm{Mpc}$, consistent with predictions from simulations and earlier SDSS-based studies. This dual detection reinforced the interpretation of splashback as a genuine feature of the matter distribution, rather than an artefact of galaxy selection or projection effects. Mpetha et al. (2025) used weak lensing profiles from the UNIONS survey to constrain cosmological parameters via the infall region structure of galaxy clusters. Their analysis yielded splashback radii of $1.59^{+0.16}_{-0.13}$, $1.30^{+0.25}_{-0.13}$, and $1.45\pm 0.11\,\mathrm{cMpc}/h$ for three independent cluster samples, and provided constraints of $\Omega_{\mathrm{m}}=0.29\pm 0.05$ and $\sigma_{8}=0.80\pm 0.04$. These results demonstrate that splashback-based lensing analyses can break degeneracies inherent in traditional abundance and shear-based cosmological tests.
However, lensing studies also highlight systematic challenges. Zhang et al. (2023) showed that neglecting the splashback feature in mass profile modelling can bias weak lensing mass estimates, particularly for group-sized halos ($M_{200m}\sim 10^{13.5}\,M_{\odot}$), by more than 0.1 dex. This bias propagates into cosmological analyses, potentially leading to underestimates of $\Omega_{\mathrm{m}}$. Umetsu & Diemer (2017) further demonstrated that the splashback feature is most pronounced when cluster profiles are scaled by $R_{200m}$, and that averaging over physical radii can smear out the steepening signal, underscoring the importance of profile scaling in lensing analyses.
Overall, lensing studies have established the splashback radius as a robust and observable feature of cluster mass profiles, with significant implications for halo boundary definitions and cosmological parameter estimation. As future surveys such as Euclid and LSST deliver deeper and wider lensing data, splashback-based analyses are poised to become a key tool in precision cosmology.
SZ-selected cluster samples offer a redshift-independent, mass-limited catalogue of clusters with minimal projection bias. Studies using clusters from Planck, SPT, and ACT have demonstrated that the splashback radius inferred from the galaxy distribution in SZ-selected samples closely matches expectations from simulations (e.g. Shin et al., 2019; Zürcher & More, 2019; Adhikari et al., 2021). In contrast, optically selected clusters—such as those in redMaPPer samples—often yield smaller splashback radii due to selection and projection effects. Because SZ selection is more tightly correlated with halo mass, it provides an important benchmark for testing the physical interpretation of the splashback feature.
Our analysis of the stellar splashback feature, traced through the total stellar density profiles, adds a new observational window on this boundary. We find that the truncation radius of the stellar orbiting component coincides with that of the dark matter and follows the same dependence on the mass accretion rate, $\Gamma$, in agreement with earlier simulation-based studies (Deason et al., 2021; Diemer, 2025). The steeper fall-off in the stellar profile makes the splashback edge more sharply defined than in lensing profiles, and may be observable through deep imaging of the ICL. These findings validate the use of projected stellar profiles as observational proxies for cluster boundaries. While subject to baryonic effects, the close alignment with dark matter splashback structures highlights the value of stellar tracers for probing cluster dynamics.
Taken together, lensing, SZ, and stellar-based methods provide a powerful set of tools for characterising the outskirts of galaxy clusters. Our results, which show that the stellar splashback radius coincides with that of the dark matter and follows the same accretion rate dependence, suggest that stellar tracers—especially the ICL—can provide a complementary and potentially sharper observational probe of cluster boundaries, provided that projection effects and substructure contamination are properly accounted for.
Each is sensitive to different physical components—total mass, thermal pressure, and stellar material—and affected by distinct systematics. The convergence of splashback measurements across these methods strengthens the interpretation of this feature as a physically meaningful halo boundary and supports its use as a probe of cluster growth and accretion history (e.g. Joshi et al., 2025).
6 Conclusions
We have used hydrodynamical simulations from The Three Hundred Project to study the splashback structure of galaxy clusters, focusing on the stellar and dark matter components. By decomposing each into orbiting and infalling material, we quantified their density profiles and characterised the splashback feature through the truncation radius $r_{\mathrm{t}}$ and slope $\beta$. Our key findings are as follows:
  1. 1.
    Coincidence of Splashback Radius: The splashback radius $r_{\rm t}$ of the stellar component coincides with that of the dark matter on a halo-by-halo basis, confirming a close dynamical coupling between stars and dark matter in the cluster outskirts. This supports earlier results from Deason et al. (2021), and extends them by demonstrating this agreement across a statistically representative sample of clusters.
  2. 2.
    Sharper Stellar Truncation: The stellar orbiting profile exhibits a steeper truncation than the dark matter, consistent with a sharper splashback feature. This is in line with the findings of Deason et al. (2021) and O’Neil et al. (2021), and suggests that stellar material - in particular the ICL - may offer a more distinct observational signature of the splashback boundary.
  3. 3.
    Accretion Rate Dependence: Both components follow the same $r_{\rm t}-\Gamma$ relation, enabling the use of stellar profiles as tracers of halo mass accretion history. This builds on the theoretical framework established by Adhikari et al. (2014), More et al. (2015), and Diemer (2017), and confirms that the splashback radius is sensitive to recent mass growth, with minimal dependence on halo mass or cosmology.
  4. 4.
    Observational Viability: Projected stellar density profiles allow recovery of the splashback radius $R_{\rm t}$ with modest scatter ($\sim 0.3R_{200m}$), demonstrating its observational viability. This complements observational detections of splashback features in galaxy density and weak lensing profiles (e.g. Chang et al. 2018b; Baxter et al. 2017; Gabriel-Silva & Sodré 2025) and supports the use of the ICL as a tracer of cluster boundaries, especially in deep imaging surveys.
  5. 5.
    Infalling and Inner Profile Correlations: The infalling and inner stellar profiles correlate with those of the dark matter, particularly in non-parametric measures such as the scale radius $r_{s}$. This suggests that stellar material can be used to probe both the outer and inner dynamical structure of clusters, consistent with results from Pizzardo et al. (2024) and O’Neil et al. (2021).
These results establish the stellar splashback feature—detectable via the ICL or integrated stellar profiles—as a robust, physically motivated tracer of cluster boundaries and growth rates. This opens a path for future studies to exploit stellar observables in constraining the dynamical state and assembly history of clusters, complementing traditional methods such as satellite kinematics and weak lensing.
As upcoming surveys like Euclid, LSST, and Roman deliver deeper imaging and improved statistics, the stellar splashback radius may become a key tool in precision cluster cosmology. For this reason, there should be a focus on developing observational pipelines for ICL-based splashback detection, quantifying systematics due to projection and substructure, and integrating splashback measurements into multi-probe cosmological frameworks (e.g. Mpetha et al., 2025; Joshi et al., 2025).
Acknowledgements
KW and CP acknowledge the support of the ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013. AK is supported by the Ministerio de Ciencia e Innovación (MICINN) under research grant PID2021-122603NB-C21 and further thanks Hüsker Dü for new day rising. WC is supported by the STFC AGP Grant ST/V000594/1, the Atracción de Talento Contract no. 2020-T1/TIC-19882 was granted by the Comunidad de Madrid in Spain, and the science research grants were from the China Manned Space Project. He also thanks the Ministerio de Ciencia e Innovación (Spain) for financial support under Project grant PID2021-122603NB-C21 and HORIZON EUROPE Marie Sklodowska-Curie Actions for supporting the LACEGAL-III project with grant number 101086388.This work has been made possible by the The Three Hundred collaboration [^1] https://www.the300-project.org. The HD simulations (7K and 15K runs) were performed on the MareNostrum Finisterrae3, and Cibeles Supercomputers through The Red Española de Supercomputación grants (AECT-2022-3-0027, AECT-2023-1-0013, AECT-2023-2-0004, AECT-2023-3-0023, AECT-2024-1-0026), on the DIaL3 – DiRAC Data Intensive service at the University of Leicester through the RAC15 grant: Seedcorn/ACTP317, and on the Niagara supercomputer at the SciNet HPC Consortium. DIaL3 is managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. This work also used the DiRAC Complexity system, operated by the University of Leicester IT Services, which forms part of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). This equipment is funded by BIS National E-Infrastructure capital grant ST/K000373/1 and STFC DiRAC Operations grant ST/K0003259/1. DiRAC is part of the National e-Infrastructure. SciNet (Loken et al., 2010) is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.
Data Availability
All data used in our analysis are from The Three Hundred collaboration.